OpenStudy (anonymous):

The length of the projection of a line on a plane is 12 cm. If the length of the line is 18 cm, find the angle it makes with the plane?

OpenStudy (anonymous):

Call the angle φ. 18 cosφ = 12 cosφ = 12/18 = 2/3 \[φ\approx48.2^{o}\]